OpenStudy (anonymous):

4^(x-2) - 3^x >=0 FInd the domain ??

OpenStudy (jango_in_dtown):

First write \[4^{x-2}\ge 3^{x}\]

OpenStudy (jango_in_dtown):

then take logarithms on both sides and you have \[(x-2)\ln 4\ge x \ln3\]

OpenStudy (jango_in_dtown):

then you have \[x (\ln4-\ln 3)\ge 2 \ln4\]

OpenStudy (jango_in_dtown):

the imply x>= ln 16/ ln(4/3) since ln 4-ln 3=ln (4/3) and 2 ln 4=ln 16

OpenStudy (jango_in_dtown):

this imply x>=9.637

OpenStudy (jango_in_dtown):

@Klevis